OpenStudy (anonymous):

For any circle the ratio of the radius to the circumference is approximately: a. 6 to 1 b. 3 to 1 c. 1 to 3 d. 1 to 6

OpenStudy (anonymous):

1 to 6

OpenStudy (dumbcow):

C = 2pi*r r/C = r/2pi*r = 1/2pi

OpenStudy (binary3i):

\[2\pi/1\]

OpenStudy (binary3i):

reverse it

OpenStudy (anonymous):

can you explain please

OpenStudy (anonymous):

So we want the ratio of the radius to circumference. radius is r and circumference is C=2 pi r radius:circumference = r:2pi*r I divide both sides by r to get = 1:2pi = 1:6.2831... So the answer will 1 to 6 as that it the closest.

OpenStudy (anonymous):

thanx zed